# LUDVIG: Learning-free Uplifting of 2D Visual features to Gaussian Splatting scenes

### v4

@@ A.4 Open-vocabulary object localization
For the open-vocabulary object localization task, multi-resolution CLIP feature maps are constructed as described in Sec. 4.2 and uplifted along with DINOv2 features using Eq. (3). The refined 3D CLIP features are then evaluated on the LERF localization and segmentation tasks as described below.
@@
We consider uplifted 3D CLIP features $f$. We follow LERF (Kerr et al., 2023) and LangSplat (Qin et al., 2024) to compute alignment scores between CLIP visual features and a text query, denoted as relevancy score, and for object localization based on these relevancy scores.
Relevancy scores.
The relevancy $r_{i,q}$ between a feature $f_{i}$ and text query $\phi_{\text{q}}$ is defined as follows:
@@
where $T$ is a temperature parameter set to 10 by  Kerr et al. (2023) and $\phi_{\text{cano}}^{k}$ is the CLIP embedding of a predefined canonical phrase chosen from “object,” “things,” “stuff,” and “texture.” Note that Qin et al. (2024) compute the relevancy scores for 2D pixels, while we directly compute them for 3D Gaussians, allowing their manipulation in 3D.
Localization.
The 3D CLIP relevancies can be projected into 2D for a given camera pose, and used for localization and segmentation for each text query. For localization, we follow Qin et al. (2024) and choose the pixel with the highest relevancy score, following a 2D smoothing using a mean filter with kernel size $K=5$ in our work.
@@ A.4.2 Object segmentation
Given a camera pose and a text query, a segmentation mask is obtained by first applying a rough thresholding over projected relevancies rescaled by their maximum value, with a fixed threshold value $\tau=0.8$, followed by automatic thresholding with Otsu’s method (Otsu et al., 1975).
Improving 2D segmentation with SAM.
For segmentation with SAM, we use the pixels with the highest relevancy scores as query points for a given camera pose and text query. More precisely, we first obtain a mask $\mathcal{M}$ by projecting and thresholding the relevancies as described above, using $\tau=0.93$. We then use the approach described in Sec. A.1, paragraph Predicting a SAM segmentation mask from a set of query points and average 20 mask predictions. We choose the top-$q$ percent of pixels as the subset of query points for SAM, where $q$ is the proportion of positive pixels in $\mathcal{M}$, hence extracting a larger subset of point prompts for larger objects. The scalar map obtained by averaging the 20 predicted masks is then automatically thresholded again using Otsu’s method (Otsu et al., 1975).
Refining 3D CLIP relevancies with graph diffusion based on DINOv2 features.
@@
We refine 3D CLIP relevancy scores using graph diffusion based on 3D DINOv2 features ($f$), as in Sec. A.3. The diffusion process runs in parallel for all text queries. For initialization, we keep positive a very restricted set of nodes with the highest relevancy, whose weight propagate to neighboring nodes with similar DINOv2 features, progressively spanning the object of interest. The diffusion process results in a better segmentation both with and without leveraging SAM. When using SAM, the set of query pixels can have a larger coverage of the object of interest without extending to other objects.
Details on graph construction and node initialization for refining 3D CLIP relevancies.
@@
The pairwise similarity function $S_{f}$ is defined as in Eq. 9 with a bandwidth value $b=0.5$. For each text query $\phi_{q}$, we define a unary regularization term $P_{q}$ using a logistic regression model $\mathcal{L}_{q}$ that predicts the probability that a DINOv2 feature $f_{i}$ belongs to the object corresponding to query $\phi_{q}$. The set of nodes $\mathcal{P}$ with positive labels is defined based on 3D CLIP relevancy scores for prompt $\phi_{q}$. More precisely, we rescale 3D relevancies to $[0,1]$ and apply Otsu’s method (Otsu et al., 1975) over relevancies above $0.5$. We use a regularization $C=0.001$ and equal class weighting for training the model. The unary regularization term $P_{q}$ is then defined as $P_{q}(f_{i})=\mathcal{L}_{q}(f_{i})^{1/b}$, with $b=0.025$ for segmentation with SAM, and $b=2$ otherwise. The initial weight vector $g_{0}$ is defined by applying two more iterations of Otsu’s method among nodes in $\mathcal{P}$ and setting to zero all relevancies below the given threshold. Restricting the set of initial points ensures diffusion only happens within the object of interest. Segmentation based on the resulting 3D relevancy scores is then performed as described in the previous paragraphs, using $\tau=0.01$ for segmentation with SAM and $\tau=0$ otherwise.
Appendix B Additional results
B.1 Runtime analyses
In this section, we detail our running times for feature uplifting and evaluation, conducted on a GPU A6000 ADA. Table 5 shows a breakdown of running times between feature uplifting (Up.) and generation (Gen.), graph diffusion and 2D segmentation for evaluation on LERF segmentation. The total reported times can be divided between pre-uplifting, uplifting and post-uplifting. In our experiments, the pre-processing and uplifting steps are independent from the downstream tasks (except for our foreground/background segmentation with SAM), and part of the graph-diffusion process is task-dependent. Below we detail our runtimes for every step and compare them to the literature.
Pre-uplifting: feature map generation.
The time this step takes (Gen. in Table 5) depends on the backbone model, on the number of training images and on the number of calls to the model per image. The total time ranges from a few seconds up to an hour for approaches such as LangSplat (Qin et al., 2024), that queries SAM over a grid of points on the image at various resolutions to generate full image segmentation masks. This process takes 24s/image on a GPU 6000 ADA and amounts to an average of 80 minutes for the evaluated scenes. In our experiments, the feature generation step takes from 1 to 5 minutes.
Uplifting.
For LUDVIG, uplifting time is linear in the number of images (given a 3D scene representation). Experimentally, it is also linear in the number of feature dimensions, taking 2ms per dimension for an image of size $724\times 986$. As reported in Table 5 (Up.), uplifting 100 images of dimension $40$ takes 9s on average. By contrast, gradient-based optimization requires approximately $n_{\text{steps}}$ times this duration, where the number of gradient steps $n_{\text{steps}}$ typically ranges from 3,000 to 30,000 for 3D feature distillation (Kerr et al., 2023; Qin et al., 2024; Zuo et al., 2024). Gradient-based optimization can still be very fast for low-dimensional features such as SAM masks (can take as little as a few seconds, as reported by SA3D-GS (Cen et al., 2023b)) or dimensionality-reduced features (LangSplat (Qin et al., 2024) trains an autoencoder to reduce the CLIP feature dimension from 512 to 3 and runs for 25 minutes). However, optimization becomes intractable for high-dimensional features such as CLIP and DINO; FMGS (Zuo et al., 2024) relies on an efficient multi-resolution hash embedding of the scene; however, their total training time still amounts to 1.4 hours.
Post-uplifting: graph diffusion.
After uplifting, LUDVIG performs graph diffusion using pairwise DINOv2 feature similarities for segmentation tasks. In Table 5, we divide runtimes in two categories:
  • Scene: operations performed once for the whole scene. This includes querying the Euclidean neighbors for each node, which is log-linear in the number of Gaussians. With 600,000 Gaussians as in our experiments, the step takes about 30s, and can be further optimized by using approximate nearest neighbor search algorithms (Wang et al., 2021). Defining $S_{f}$ based on DINOv2 features is also independent from the downstream task.
Post-uplifting: segmentation.
Our evaluation on LERF involves 2D segmentation with SAM based on 3D relevancy scores. The runtime depends on the number of test images and on the total number of 2D text queries, as it involves one call to the SAM backbone per test image, and multiple calls to the SAM prediction head per text query, as detailed in Appendix A.4. Our total inference time per scene is of $8s$ on average, against $0.8s$ when not using SAM. In contrast, Langsplat does not rely on SAM at inference time, but relies on a computationally expensive feature map generation process, with more than 1 hour runtime.
@@ B.1 Runtime analyses
Table 5: Runtimes for evaluation on LERF Segmentation (Kerr et al., 2023; Qin et al., 2024). The last column (Total) reports total time, which breaks down between i) feature map generation (Gen.) and uplifting (Up.) for all training images; ii) graph diffusion, divided between scene-specific (querying neighbors, defining $S_{f}$) and prompt-specific (defining $P$, running diffusion) operations for all text queries; iii) 2D segmentation with/without SAM for all text queries across test images. We also report the number of training and test images and the number of text queries across test images.
@@
Table 6: Segmentation (IoU) on SPIn-NeRF (Mirzaei et al., 2023) with DINOv2, SAM and SAM2.
@@
Table 7: Segmentation (IoU) on NVOS (Ren et al., 2022) with DINOv2, SAM and SAM2.
@@
Table 8: Segmentation (IoU) on SPIn-NeRF (Mirzaei et al., 2023). We compare purely geometrical reference mask uplifting and reprojection, single-view prediction, feature/mask uplifting, and graph diffusion leveraging DINOv2 or SAM2.
B.2 Per-scene foreground/background segmentation results
@@
In this section, we present per-scene segmentation results on NVOS and SPIn-NeRF in Tables 67 and 8, along with an extended analysis of these results.
Segmentation on SPIn-NeRF.
@@
We report our segmentation results for the SPin-NeRF dataset (Mirzaei et al., 2023) in Table 6. Our results are comparable to the state of the art while not relying on optimization-based approaches. Surprisingly, our segmentation with DINOv2 using graph diffusion also gives results on par with models leveraging SAM masks. Our lower segmentation results compared to OmniSeg’s can be partly attributed to poor Gaussian Splatting reconstruction of highly specular scenes such as the Fork, in which semi-transparent Gaussians floating over the object try to represent reflections or surface effects that are difficult to capture with standard rasterization techniques (Jiang et al., 2024).
Segmentation on NVOS.
@@
We report our segmentation results for the NVOS dataset (Ren et al., 2022) in Table 7. Our results are comparable to those obtained by prior work. Again, DINOv2 performs surprisingly well while not having been trained on billions of labeled images like SAM. Compared to SAM, DINOv2 better captures complex objects, but sometimes also captures some background noise. This can be seen in Appendix Figure 5 with the example of Trex: while SAM misses out the end of the tail as well as the end of the ribs, DINOv2 captures the whole Trex, but also captures part of the stairs behind. Visualisations of Orchids in Appendix Figure 5 also explain the lower performance of SAM on this scene: the two orchids SAM is missing are not covered by the positive scribbles, which makes the task ambiguous.
Ablation study.
@@
In Table 8, we compare our segmentation protocol using DINOv2 and SAM2 to multiple simpler variants. More precisely, we evaluate i) a purely geometrical variant that does not use SAM2 or DINOv2, ii) single-view segmentation in 2D based on SAM2 or DINOv2 2D predictions, iii) uplifting DINOv2 features or SAM2 masks into 3D then rendering them for segmentation, as described in Sec. A.1 and A.2, and iv) segmenting using graph diffusion over DINOv2 3D feature similarities.
@@
The purely geometrical approach works well on the forward-facing LLFF scenes (Orchids to Fortress). In these scenes, the reference mask is accurately uplifted and reprojected as the viewing direction changes only a little between each frame. However, it fails on the 360-degree scenes (Fork, Pinecone, Truck, Lego). This points to a suboptimal 3D reconstruction of the scene, likely due to overfitting on the limited numbers of available training views (Chung et al., 2024).
@@
The single-view variants use a similar process for constructing the features and using them for segmentation as in Sec. A.1 and A.2 but without uplifting and rendering. It improves from a purely geometrical approach and performs reasonably well on average, the foreground being well isolated from the rest of the scene. However, as illustrated in Figure 2, the semantic features are at a much lower resolution than those resulting from 3D uplifting, leading to a coarser segmentation.
@@
3D uplifting considerably boosts results compared to single-view approaches. However, performing segmentation in 2D based on the uplifted DINOv2 features does not benefit from the 3D spatial information and typically fails on the 360-degree scenes (Pinecone, Truck and Lego) which have higher variability between frames from different views. Introducing 3D spatial information through 3D graph diffusion results in a boosted performance on these scenes.
Appendix C Additional visualizations
C.1 Segmentation tasks
Segmentation on NVOS.
@@ C.1 Segmentation tasks
Figure 5 shows our segmentation masks from SAM and DINOv2 for the three most challenging scenes of the NVOS dataset: Fern, Orchids and Trex.
@@
Figure 5: Segmentation results on NVOS (Ren et al., 2022) with DINOv2 and SAM.
Diffusion process.
Figure 6 illustrates different steps of the diffusion process for Fern, Leaves, Flower and Trex from the NVOS (Ren et al., 2022) dataset. Starting from the reference scribbles, the diffusion rapidly spreads through the large neighboring Gaussians. Covering the entire object takes more time for complex structures such as Fern, or for masks with disconnected components such as Orchids. As illustrated in the case of Flower, the last diffusion steps allow spreading to the smaller Gaussians on the flowers’ edges, yielding a refined segmentation mask. For Trex, the parts being reached the latest are the head and tail. Their features are further away from the reference features (defined as the average feature over 3D reference scribbles), and therefore the regularization for diffusion is stronger in these regions. Overall once the object has been fully covered, the regularization is very effective at preventing leakage, which allows diffusion to run for an arbitrary number of steps.
Figure 6: Illustration of the graph diffusion process. 2D projections of i) first three PCA components of DINOv2 3D features, ii) unary regularization term (red), iii) weight vector $g_{t}$ at timesteps $t\in\{0,3,5,10,100\}$, iv) RGB segmentation obtained using a mask based on the 2D projection of $g_{100}$.
Scene editing.
Figure 7 shows comparative visualizations of scene editing with N3F (Tschernezki et al., 2022) and LUDVIG. For rendering the edited RGB image, N3F sets to zero the occupancy for all 3D points belonging to the object. For LUDVIG, we remove all Gaussians pertaining to the 3D semantic mask resulting from graph diffusion. We observe that the regions behind to segmented object are much smoother for LUDVIG than for N3F. Regions unseen from any viewpoint are black for LUDVIG (no gaussians) and result in a background partially hallucinated by NeRF for N3F.
@@
(b) N3F (Tschernezki et al., 2022)
(c) LUDVIG
Figure 7: Scene editing. 3D segmentation, removal and rendering for LUDVIG and N3F (Tschernezki et al., 2022). For N3F, figures are sourced from (Tschernezki et al., 2022).
C.2 Visual comparisons of uplifted features
Figure 8 show a comparison of LUDVIG’s 3D DINOv2 features with learned 3D DINO features of N3D (Tschernezki et al., 2022). Their figures are taken from their work. The notable differences are a more fine-grained reconstruction of the background for the trex and horns, and overall smoother features across all scenes.
Figure 8: Comparison between LUDVIG’s uplifted DINOv2 features (bottom) and N3F’s (Tschernezki et al., 2022) learned DINO features (top). For N3F, figures are sourced from  (Tschernezki et al., 2022).
C.3 Comparison to GaussianEditor’s uplifting.
Our aggregation procedure in Eq. (5), illustrated in Figure 1, bears similarity with the one from Chen et al. (2024) for uplifting 2D binary masks to a 3D Gaussian splatting scene. In their method, uplifted masks are thresholded to create 3D binary masks that are used for semantic tracing. Specifically, they rely on rough 3D segmentation masks to selectively optimize Gaussians that are relevant for an editing task. Unlike in Eq. (3) and (5), Chen et al. (2024) propose to normalize their uplifted masks based on the total count of view/pixel pairs $(d,p)$ contributing to the mask of a Gaussian $i$, i.e. $\sum_{d,p\in\mathcal{S}_{i}}1$, without taking the rendering weight $w_{i}(d,p)$ into account. Consequently, the uplifted features tend to have larger values for large, opaque Gaussians. Figure 9 shows a qualitative comparison between 3D DINOv2 features obtained using the aggregation proposed by Chen et al. (2024) and our approach. The aggregation by Chen et al. (2024) fails to assign the right semantics to large gaussians, which is particularly visible in scenes with high specularity such as Room. This showcases the importance of defining 3D features as convex combinations of 2D pixel features.

### v5

@@ A.4 Open-vocabulary object localization
For the open-vocabulary object localization task, multi-resolution CLIP feature maps are constructed as described in Sec. 4.2 and uplifted along with DINOv2 features using Eq. (3) from the main paper. The refined 3D CLIP features are then evaluated on the LERF localization and segmentation tasks as described below.
@@
We consider uplifted 3D CLIP features $f$. We follow LERF [19] and LangSplat [35] to compute alignment scores between CLIP visual features and a text query, denoted as relevancy score, and for object localization based on these relevancy scores.
Relevancy scores. The relevancy $r_{i,q}$ between a feature $f_{i}$ and text query $\phi_{\text{q}}$ is defined as follows:
@@
where $T$ is a temperature parameter set to 10 by [19] and $\phi_{\text{cano}}^{k}$ is the CLIP embedding of a predefined canonical phrase chosen from “object,” “things,” “stuff,” and “texture.” Note that [35] compute the relevancy scores for 2D pixels, while we directly compute them for 3D Gaussians, allowing their manipulation in 3D.
Localization. The 3D CLIP relevancies can be projected into 2D for a given camera pose, and used for localization and segmentation for each text query. For localization, we follow [35] and choose the pixel with the highest relevancy score, following a 2D smoothing using a mean filter with kernel size $K=5$ in our work.
@@ A.4.2 Object segmentation
Segmentation from 3D relevancies. Given a camera pose and a text query, a segmentation mask is obtained by first applying a rough thresholding over projected relevancies rescaled by their maximum value, with a fixed threshold value $\tau=0.8$, followed by automatic thresholding with Otsu’s method [34].
[…]
@@
Refining 3D CLIP relevancies with graph diffusion based on DINOv2 features. We refine 3D CLIP relevancy scores using graph diffusion based on 3D DINOv2 features ($f$), as in Sec. A.3. The diffusion process runs in parallel for all text queries. For initialization, we keep positive a very restricted set of nodes with the highest relevancy, whose weight propagate to neighboring nodes with similar DINOv2 features, progressively spanning the object of interest. The diffusion process results in a better segmentation both with and without leveraging SAM. When using SAM, the set of query pixels can have a larger coverage of the object of interest without extending to other objects.
@@
Details on graph construction and node initialization for refining 3D CLIP relevancies. The pairwise similarity function $S_{f}$ is defined as in Eq. 9 with a bandwidth value $b=0.5$. For each text query $\phi_{q}$, we define a unary regularization term $P_{q}$ using a logistic regression model $\mathcal{L}_{q}$ that predicts the probability that a DINOv2 feature $f_{i}$ belongs to the object corresponding to query $\phi_{q}$. The set of nodes $\mathcal{P}$ with positive labels is defined based on 3D CLIP relevancy scores for prompt $\phi_{q}$. More precisely, we rescale 3D relevancies to [0,1][0,1][ 0 , 1 ] and apply Otsu’s method [34] over relevancies above 0.50.50.5. We use a regularization $C=0.001$ and equal class weighting for training the model. The unary regularization term $P_{q}$ is then defined as $P_{q}(f_{i})=\mathcal{L}_{q}(f_{i})^{1/b}$, with $b=0.025$ for segmentation with SAM, and $b=2$ otherwise. The initial weight vector $g_{0}$ is defined by applying two more iterations of Otsu’s method among nodes in $\mathcal{P}$ and setting to zero all relevancies below the given threshold. Restricting the set of initial points ensures diffusion only happens within the object of interest. Segmentation based on the resulting 3D relevancy scores is then performed as described in the previous paragraphs, using $\tau=0.01$ for segmentation with SAM and $\tau=0$ otherwise.
@@
Table A: Runtimes for evaluation on LERF Segmentation [19, 35]. The last column (Total) reports total time, which breaks down between i) feature map generation (Gen.) and uplifting (Up.) for all training images; ii) graph diffusion, divided between scene-specific (querying neighbors, defining $S_{f}$) and prompt-specific (defining $P$, running diffusion) operations for all text queries; iii) 2D segmentation with/without SAM for all text queries across test images. We also report the number of training and test images and the number of text queries across test images.
SA-TensoRF SA-GS OmniSeg3D SAGA LUDVIG
SAM DINOv2
2D feature extraction (sec.) 30 30 90022footnotemark: 2 90011footnotemark: 1 20 30
3D feature training (sec.) 15 4 900 1500 4 40+2
(a) Multi-view segmentation (NVOS). Italic denote inference-time operations (after 2D scribbles are given). OmniSeg3D and SAGA avoid inference overhead but require longer feature extraction and training.
LERF LangSplat OpenGaussian LUDVIG
w/o diff. w/ diff.
2D feature extraction (min) -  50505011footnotemark: 1  50505011footnotemark: 1 4.5 4.5
3D feature training (min) 25 25 40 2 2.5
Inference (s/query) 30.9 0.26 0.2$\sim$ 0.2∼ 0.2 0.3 1.3
(b) Open-vocabulary segmentation (Ramen scene) .
Table B: Training and inference times. All methods build on top of a GS pretraining stage (10$\sim$ 10∼ 10min). Reported values are paper-sourced and indicative, as hardware setups may differ.
Appendix B Runtime analysis
B.1 Runtimes for LUDVIG
In this section, we detail our running times for feature uplifting and evaluation, conducted on a GPU RTX 6000 ADA. Table A shows a breakdown of running times between feature uplifting (Up.) and generation (Gen.), graph diffusion and 2D segmentation for evaluation on LERF segmentation. The total reported times can be divided between pre-uplifting, uplifting and post-uplifting. In our experiments, the pre-processing and uplifting steps are independent from the downstream tasks (except for our foreground/background segmentation with SAM), and part of the graph-diffusion process is task-dependent. Below we detail our runtimes for every step and compare them to the literature.
Pre-uplifting: feature map generation. The time this step takes (Gen. in Table A) depends on the backbone model, on the number of images and on the number of calls to the model per image. The total time ranges from a few seconds up to an hour for approaches such as LangSplat [35], that queries SAM over a grid of points on the image at various resolutions to generate full image segmentation masks. [^3] This process takes 24s/image on a GPU 6000 ADA and amounts to an average of 80 minutes for the evaluated scenes. In our experiments, the feature generation step takes from 1 to 5 minutes, except for Sec. 5.4 where we uplift semantic maps generated with LangSplat’s approach.
Uplifting. For LUDVIG, uplifting time is linear in the number of images (given a 3D scene representation). Experimentally, it is also linear in the number of feature dimensions, taking 2ms per dimension for an image of size 724×986724$\times$ 986724 × 986. As reported in Table A (Up.), uplifting 100 images of dimension 404040 takes 9s on average. By contrast, gradient-based optimization requires approximately $n_{\text{steps}}$ times this duration, where the number of gradient steps $n_{\text{steps}}$ typically ranges from 3,000 to 30,000 for 3D feature distillation [19, 35, 56]. Gradient-based optimization can still be very fast for low-dimensional features such as SAM masks (can take as little as a few seconds, as reported by SA3D-GS [3]) or dimensionality-reduced features (LangSplat [35] trains an autoencoder to reduce the CLIP feature dimension from 512 to 3 and runs for 25 minutes). However, optimization becomes intractable for high-dimensional features such as CLIP and DINO; FMGS [56] relies on an efficient multi-resolution hash embedding of the scene; however, their total training time still amounts to 1.4 hours.
Post-uplifting: graph diffusion. After uplifting, LUDVIG performs graph diffusion using pairwise DINOv2 feature similarities for the segmentation tasks in Sec. 5.2 and 5.3. In Table A, we divide runtimes in two categories:
  • Scene: operations performed once for the whole scene. This includes querying the Euclidean neighbors for each node, which is log-linear in the number of Gaussians. With 600,000 Gaussians as in our experiments, the step takes about 30s, and can be further optimized by using approximate nearest neighbor search algorithms [45]. Defining $S_{f}$ based on DINOv2 features is also independent from the downstream task.
[…]
B.2 Runtime comparisons
Table B reports runtime comparisons for the multi-view (Sec. 5.2) and open-vocabulary (Sec. 5.3) segmentation tasks. Compared to approaches relying on SAM’s automatic mask generation, such as OmniSeg3D [52], SAGA [5], LangSplat [35], and OpenGaussian [46], our method offers significantly faster feature generation and uplifting, at the cost of a slightly higher inference time. In particular, the graph diffusion step adds approximately 1 second per query at inference (reported as 14 seconds for Ramen’s 14 queries in supplementary Table A).
Our experiments in Sec. 5.4 show that our uplifting process can serve as a drop-in replacement for the training phase in methods optimized for fast inference, such as OpenGaussian [46], which learns from 2D feature maps generated using LangSplat’s approach. In these experiments, we adopt the same 2D feature map generation and evaluation protocols as OpenGaussian, replacing only their quantization-based training phase with our fast feature aggregation. This results in significantly faster uplifting while achieving notable accuracy gains.
@@ B.2 Runtime comparisons
Table C: Segmentation (IoU) on SPIn-NeRF [32] with DINOv2, SAM and SAM2.
@@
Table D: Segmentation (IoU) on NVOS [39] with DINOv2, SAM and SAM2.
@@
Table E: Segmentation (IoU) on SPIn-NeRF [32]. We compare purely geometrical reference mask uplifting and reprojection, single-view prediction, feature/mask uplifting, and graph diffusion leveraging DINOv2 or SAM2.
Appendix C Additional results
C.1 Per-scene multi-view segmentation results
@@
In this section, we present per-scene segmentation results on NVOS and SPIn-NeRF in Tables CD and E, along with an extended analysis of these results.
@@
Segmentation on SPIn-NeRF. We report our segmentation results for the SPin-NeRF dataset [32] in Table C. Our results are comparable to the state of the art while not relying on optimization-based approaches. Surprisingly, our segmentation with DINOv2 using graph diffusion also gives results on par with models leveraging SAM masks. Our lower segmentation results compared to OmniSeg’s can be partly attributed to poor Gaussian Splatting reconstruction of highly specular scenes such as the Fork, in which semi-transparent Gaussians floating over the object try to represent reflections or surface effects that are difficult to capture with standard rasterization techniques [16].
@@
Segmentation on NVOS. We report our segmentation results for the NVOS dataset [39] in Table D. Our results are comparable to those obtained by prior work. Again, DINOv2 performs surprisingly well while not having been trained on billions of labeled images like SAM. Compared to SAM, DINOv2 better captures complex objects, but sometimes also captures some background noise. This can be seen in Appendix Fig. B with the example of Trex: while SAM misses out the end of the tail as well as the end of the ribs, DINOv2 captures the whole Trex, but also captures part of the stairs behind. Visualisations of Orchids in Appendix Fig. B also explain the lower performance of SAM on this scene: the two orchids SAM is missing are not covered by the positive scribbles, which makes the task ambiguous.
@@
Ablation study. In Table E, we compare our segmentation protocol using DINOv2 and SAM2 to multiple simpler variants. More precisely, we evaluate i) a purely geometrical variant that does not use SAM2 or DINOv2, ii) single-view segmentation in 2D based on SAM2 or DINOv2 2D predictions, iii) uplifting DINOv2 features or SAM2 masks into 3D then rendering them for segmentation, as described in Sec. A.1 and A.2, and iv) segmenting using graph diffusion over DINOv2 3D feature similarities.
@@
The purely geometrical approach works well on the forward-facing LLFF scenes (Orchids to Fortress). In these scenes, the reference mask is accurately uplifted and reprojected as the viewing direction changes only a little between each frame. However, it fails on the 360-degree scenes (Fork, Pinecone, Truck, Lego). This points to a suboptimal 3D reconstruction of the scene, likely due to overfitting on the limited numbers of available views [7].
@@
The single-view variants use a similar process for constructing the features and using them for segmentation as in Sec. A.1 and A.2 but without uplifting and rendering. It improves from a purely geometrical approach and performs reasonably well on average, the foreground being well isolated from the rest of the scene. However, as illustrated in Fig. 3, the semantic features are at a much lower resolution than those resulting from 3D uplifting, leading to a coarser segmentation.
@@
3D uplifting considerably boosts results compared to single-view approaches. However, performing segmentation in 2D based on the uplifted DINOv2 features does not benefit from the 3D spatial information and typically fails on the 360-degree scenes (Pinecone, Truck and Lego) which have higher variability across different views. Introducing 3D spatial information through 3D graph diffusion results in a boosted performance on these scenes.
Appendix D Additional visualizations
D.1 Segmentation tasks
@@
Segmentation on NVOS. Fig. B shows our segmentation masks from SAM and DINOv2 for the three most challenging scenes of the NVOS dataset: Fern, Orchids and Trex.
@@
Figure B: Segmentation results on NVOS [39] with DINOv2 and SAM.
Diffusion process. Fig. C illustrates different steps of the diffusion process for Fern, Leaves, Flower and Trex from the NVOS [39] dataset. Starting from the reference scribbles, the diffusion rapidly spreads through the large neighboring Gaussians. Covering the entire object takes more time for complex structures such as Fern, or for masks with disconnected components such as Orchids. As illustrated in the case of Flower, the last diffusion steps allow spreading to the smaller Gaussians on the flowers’ edges, yielding a refined segmentation mask. For Trex, the parts being reached the latest are the head and tail. Their features are further away from the reference features (defined as the average feature over 3D reference scribbles), and therefore the regularization for diffusion is stronger in these regions. Overall once the object has been fully covered, the regularization is very effective at preventing leakage, which allows diffusion to run for an arbitrary number of steps.
Figure C: Illustration of the graph diffusion process. 2D projections of i) first three PCA components of DINOv2 3D features, ii) unary regularization term (red), iii) weight vector $g_{t}$ at timesteps $t\in\{0,3,5,10,100\}$, iv) RGB segmentation obtained using a mask based on the 2D projection of $g_{100}$.
Object removal. Fig. D shows comparative visualizations of object removal with N3F [43] and LUDVIG. For rendering the edited RGB image, N3F sets to zero the occupancy for all 3D points belonging to the object. For LUDVIG, we remove all Gaussians pertaining to the 3D semantic mask resulting from graph diffusion. We observe that the regions behind to segmented object are much smoother for LUDVIG than for N3F. Regions unseen from any viewpoint are black for LUDVIG (no gaussians) and result in a background partially hallucinated by NeRF for N3F.
@@
(b) N3F [43]
(c) LUDVIG (ours)
Figure D: Object removal. 3D segmentation, removal and rendering for LUDVIG and N3F [43]. For N3F, figures are taken from [43].
D.2 Visual comparisons of uplifted features
Fig. E show a comparison of LUDVIG’s 3D DINOv2 features with learned 3D DINO features of N3D [43]. Their figures are taken from their work. Compared to N3D, LUDVIG produces a more fine-grained reconstruction of the background (notably in the trex and horns scenes) and smoother features across all scenes.
Figure E: Comparison between LUDVIG’s uplifted DINOv2 features (bottom) and N3F’s [43] learned DINO features (top). For N3F, figures are taken from  [43].
D.3 3D features for ScanNet semantic segmentation
Fig. F presents visualizations of 3D semantic features obtained by uplifting 2D feature maps generated by OpenGaussian [46]. Following the approach of LangSplat [35], the 2D features are computed by assigning object-level CLIP embeddings to segmentation masks produced by SAM in everything mode.
Surprisingly, despite the very constrained training conditions used during Gaussian Splatting reconstruction (frozen positions and disabled densification process), the uplifting still yields coherent and meaningful 3D semantic features.
Figure F: PCA of uplifted semantic features on ScanNetv2, obtained by assigning object-level CLIP features to SAM segments.
D.4 Comparison to GaussianEditor’s uplifting
Our aggregation procedure in Eq. (3) from the main paper, illustrated in Fig. 2, bears similarity with the one from [6] for uplifting 2D binary masks to a 3D Gaussian splatting scene. In their method, uplifted masks are thresholded to create 3D binary masks that are used for semantic tracing. Specifically, they rely on rough 3D segmentation masks to selectively optimize Gaussians that are relevant for an editing task. Unlike in Eq. (3) and (5) from the main paper, [6] propose to normalize their uplifted masks based on the total count of view/pixel pairs $(d,p)$ contributing to the mask of a Gaussian $i$, i.e. $\sum_{d,p\in\mathcal{S}_{i}}1$, without taking the rendering weight $w_{i}(d,p)$ into account. Consequently, the uplifted features tend to have larger values for large, opaque Gaussians. Fig. G shows a qualitative comparison between 3D DINOv2 features obtained using the aggregation proposed by [6] and our approach. The aggregation by [6] fails to assign the right semantics to large gaussians, which is particularly visible in scenes with high specularity such as Room. This showcases the importance of defining 3D features as convex combinations of 2D pixel features.